# Predictive Coding Networks and Inference Learning: Tutorial and Survey

### v1

@@ 2.5.2. Unsupervised Learning
Unsupervised learning is concerned with finding patterns and structure in unlabeled data (bishop_pattern_2006, ). This means a change in the problem setup of section 2: we no longer have a dataset $\{\bm{x}^{(n)},\bm{y}^{(n)}\}_{n=1}^{N}$, but only $\{\bm{x}^{(n)}\}_{n=1}^{N}$. For traditional ANNs, an FNN may be adapted for unsupervised learning by using a special autoencoder architecture, encoding input data into a lower-dimensional representation followed by a decoder which reconstructs the original input. PCNs do not require a special architecture of this sort, but only change the direction of prediction to $\bm{\mu}^{\ell}=f(\bm{w}^{\ell+1}\bm{a}^{\ell+1})$ i.e. the use of a generative PCN. Indeed, PCNs were originally conceived in this way (rao_predictive_1999, ) (also see section 3).
Training Procedure. Again, one needs to only make a minor change to the approach described above. Simply keep the final layer $\bm{a}^{L}$ unclamped during training, keeping data $\bm{x}^{(n)}$ clamped to $\bm{a}^{0}$, and run IL. The model functions like an autoencoder: higher layers take the role of the latent space, which during training encodes increasingly abstract features of the data clamped to the lowest layer. As such, the PCN takes the role of both encoder and decoder.
Testing Procedure. During testing, one can sample the latent space as a decoder using a noise vector and ancestral sampling. Seeing the PCNs as a hierarchical probabilistic model (cf. section 3), each layer has a (Gaussian) conditional probability distribution. The idea of ancestral sampling is to generate a sample from the root variable(s) (the top layer) and then sample from the subsequent conditional distributions based on this (prince_understanding_2023, ). Finally, a synthetic datapoint at the bottom is obtained (ororbia_neural_2022, ).
@@ 2.5.3. Hierarchical PCNs
Discriminative and generative PCNs can together be called hierarchical PCNs, defined by the local prediction $\bm{\mu}^{\ell}=f(\bm{w}^{\ell\pm 1}\bm{a}^{\ell\pm 1})$. The supervised and unsupervised learning that can be done with generative PCNs can then be called the training modes of PCNs. These are illustrated in fig. 7. Intuitively, we can see the difference between discriminative and generative PCNs in the direction of predictions: in the discriminative model, predictions flow from data to labels, while errors flow from labels to data. In generative models, this is reversed. This is also visualized in fig. 7.
Figure 7. Overview of hierarchical PCNs. Model type (discriminative vs. generative), learning mode (supervised vs. unsupervised), and direction of predictions and errors are shown. Cf. section 2.2.1 for an explanation of our convention. For clarity, one neuron per layer is shown, and bias has been left out.
It should be noted that testing can also be ‘reversed’ in hierarchical PCNs, which we call backwards testing. If one has trained the discriminative PCN, one can clamp labels and find the minimum energy with iterative inference, producing a synthetic image at the bottom. Conversely, a trained generative PCN can classify images by iterative inference, using a clamped image instead of a clamped label. Thus, confusingly, discriminative PCNs can be used for generative tasks, and generative PCNs can be used for discriminative tasks. However, in practice, neither works better than their forward-tested counterpart, meaning generative PCN are the natural choice for generation, as discriminative PCNs are for classification. This justifies the naming convention used here and elsewhere in the literature.
@@ 2.5.4. Empirical results
Compared to discriminative PCNs, generative PCNs remain underexplored the literature. This is perhaps surprising considering the original conception of PC as an unsupervised generative model. At the same time, as was mentioned, metrics for what counts as a ‘good’ generative model are wide and varied (prince_understanding_2023, ), making it a more involved field of study in general. Indeed, the works in the literature are disparate and focus on different aspects.
Although not used for machine learning tasks, we mention (rao_predictive_1999, ) here as the first generative PCN trained in unsupervised mode. Then, (salvatori_associative_2021, ) used generative PCN as an associative memory model. It was tested on the task of reconstructing from corrupted and partial data, using up to 500 datapoints. They find good performance, outperformining autoencoders, and both Hopfield networks (hopfield_neural_1982, ) and modern Hopfield networks (krotov_dense_2016, ) in most cases. Another set of unsupervised tasks was considered by (ororbia_neural_2022, ), and extended in (ororbia_convolutional_2023, ). This work used a set of models dubbed neural generative coding (NGC). This framework is equal in spirit to PCNs as presented in section 2, but differs in notation/terminology as well as some structural aspects. The authors study the network’s performance on reconstruction & sampling ability (as measured by the likelihood) on black and white images. They find that NGC is competitive with VAEs and GANs on these tasks, and works well on downstream classification.
Next, a smaller study (kinghorn_preventing_2023, ) considers the performance of generative PCNs on standard classification tasks, training the model in supervised mode with backwards testing. Similar to (alonso_understanding_2024, ), it is observed that the weights of different layers are updated at different rates, which causes model accuracy to worsen after it has peaked. They propose that regularizing weights remedies this pathology.
The recent work of (zahid_curvature-sensitive_2023, ) considers the sampling ability of generative PCNs trained in supervised mode with a modified training objective. They find that their objective improves both likelihood and sample quality, but suffers from increased computation time with results not matching performance of VAE counterparts. An improvement was presented in (zahid_sample_2024, ), where the inference phase is combined with Langevin sampling, giving results that match or exceed VAE performance on metrics such as FID (Fréchet Inception Distance), diversity, and coverage.
@@ 2.6. PC Graphs
The previous section showed how the structure of discriminative PCNs, could be extended by changing the definition of local prediction from $\bm{\mu}^{\ell}=f(\bm{w}^{\ell-1}\bm{a}^{\ell-1})$ to $\bm{\mu}^{\ell}=f(\bm{w}^{\ell+1}\bm{a}^{\ell+1})$. This can be taken one step further: PCNs can be naturally generalized to arbitrary graphs, called PC graphs by (salvatori_learning_2022, ). [^8] Note that PC graphs are distinct from graph neural networks (GNNs), which are used for graph-structured data (prince_understanding_2023, ). These are trained using IL, but dispense with the hierarchical structure of layers. These networks can be understood as a superset of both discriminative and generative PCNs, and can flexibly be used for a variety of tasks.
@@ 2.6.1. Structure
where $w_{ij}\in\mathbb{R}$, and the sum is over all the other nodes, meaning self-connections are left out ($w_{ii}=0$), as in (salvatori_learning_2022, ). This defines a fully-connected PC graph. If we choose to include self-connections one simply takes the sum over all values of $i$. In vector notation, one can write $\bm{\mu}=f(\bm{wa})$.
A small (fully-connected) PC graph is illustrated in fig. 8, where each activation node $a_{i}$ and error node $\epsilon_{i}$ has been grouped in a vertex $v_{i}$. Importantly, graphs with different connectivity/topology can be obtained by multiplying the weight matrix with a mask or adjacency matrix. In this way, one obtains the architectures discussed in earlier sections: the discriminative PCN, as well as the generative PCN, depending on which weights are masked. This is visualized in fig. 9.
@@
Figure 9. Mask/adjacency matrices for a 4-node PC graph. Different matrices lead to different architectures, like the discriminative PCN, generative PCN, or an arbitrary topology. Larger networks are obtained in the same way. See also fig. 4 in (salvatori_learning_2022, ).
@@ 2.6.2. Training Procedure
Inference $\Delta\bm{a}^{\ell}=-\gamma\left(\bm{\epsilon}^{\ell}-\left(\bm{w}^{\ell}\right)^{T}\bm{\epsilon}^{\ell\pm 1}\odot f^{\prime}\left(\bm{w}^{\ell}\bm{a}^{\ell}\right)\right)$ $\Delta\bm{a}=-\gamma\left(\bm{\epsilon}-\bm{w}^{T}\bm{\epsilon}\odot f^{\prime}\left(\bm{w}\bm{a}\right)\right)$
@@ 2.6.4. Empirical results
PC graphs have only been used in (salvatori_learning_2022, ), with models trained using supervised mode only. They train models for classification, generation, reconstruction, denoising, and associative memory tasks (using different adjacency matrices). An interesting result is that for classification using a fully connected model, PC graphs perform much better than other fully connected architectures such as Boltzmann machines and Hopfield networks, up to 30% better on MNIST. At the same time, the fully connected model does not perform comparably to hierarchical networks trained with either BP or IL. This might be expected, considering that depth is suggested to be critical for performance in neural network learning (prince_understanding_2023, ). At the same time, such statements are based mostly on empirical evidence, and lack a definitive theoretical basis.
For the other tasks considered, only proofs of concepts are provided to demonstrate their flexibility. Unsupervised learning is also not considered.
@@ 2.7. Generalized ANNs
Thus, it becomes clear that formally, PCNs and PCGs can be seen as types of ‘generalized ANNs’ that go beyond hierarchical structures, by virtue of the use of IL as opposed to BP. This is very interesting, for at least two reasons. First, as observed by (salvatori_learning_2022, ), PC graphs allow one to train non-hierarchical structures with a brain-like topology. Speculatively, such networks could, if better understood, share some of the advantages that biological brains have over ANNs, such as vastly superior energy efficiency and parallelization. Second, from a very general perspective, topological considerations have strongly contributed to several advances in machine learning in the past. A prominent example is that of residual networks and skip connections (zagoruyko_wide_2017, ), which have allowed training of much deeper networks, and can improve performance on a variety of tasks (prince_understanding_2023, ). In a sense, the very notion of depth in deep learning is a topological feature. As such, we consider this an important avenue for further work.
@@ 3. PCNs as Probabilistic Latent Variable Models
Section 2 introduced the PCN as a type of generalized ANN. In this section we discuss a second, complementary, perspective: PCN as a probabilistic latent variable model. Conceptualizing PCNs in this way allows a principled derivation of equations in the previous section, and brings to light connections to other methods well-known in machine learning (e.g., energy-based models and linear factor models). It also reveals a number of assumptions and modeling choices, providing a basis for improvements and possible future developments.
We start our step-by-step derivation from maximum likelihood estimation in probabilistic (Bayesian) models, discussing how expectation maximization (EM) implements maximum likelihood for models with latent variables, and how different generative models lead to PC as formulated by (rao_predictive_1999, ) or to multi-layer PC as described in the previous section.
@@ A Brief History of PCNs
The original model of Rao & Ballard (rao_predictive_1999, ) was conceived as a hierarchical model with Gaussian relations between layers. This formulation contained the ingredients typically associated with PC, chief among them the minimization of prediction errors in a hierarchical fashion. Work by Friston (friston_learning_2003, ; friston_free_2006, ; friston_hierarchical_2008, ) showed how the computational steps in PC (minimization of the energy with respect to activations and weights) are an instance of EM, also understood as a variational inference procedure using a variational free energy (i.e., the evidence lower bound) as the objective function. This perspective explains why EM maximizes the likelihood of the data, and can be used to derive generalized objective functions (see appendix A).
@@ 3.1. Expectation Maximization
(17) $\operatorname{NLL}(\theta)=-\ln p_{\theta}(\bm{x})$
For some joint distribution $p_{\theta}(\bm{x},\bm{z})$, also called the generative model, it can be shown that the following two steps will minimize the NLL (see appendix A):
(18) $\widetilde{\bm{z}}$ $=\underset{\bm{z}}{\operatorname{argmax}}\,p_{\theta}(\bm{x},\bm{z})$ (E-step)
@@
Here, $\ln p_{\theta}(\bm{x},\bm{z})$ is also called complete data log-likelihood, since it represents the likelihood function calculated using the full set of observed and latent variables – to distinguish it from the likelihood calculated only with observed data (bishop_pattern_2006, ). For convenience, we will label the negative complete data log-likelihood as $E$, following:
@@
which we will choose to minimize, as opposed to maximizing $\ln p_{\theta}(\bm{x},\bm{z})$. This, with a change of notation, gives IL, exactly as presented in section 2. It has furthermore been shown (cf. appendix A, ref. (neal_view_1998, )) that partial E-steps and M-steps will also minimize the NLL:
(21) $\Delta\bm{z}$ $\propto-\frac{\partial E}{\partial\bm{z}}$ (partial E-step)
(22) $\Delta{\theta}$ $\propto-\frac{\partial E}{\partial{\theta}}$ (partial M-step)
which is indeed how IL is implemented in practice.
@@ 3.2.1. PCNs of Rao & Ballard
and likewise for the prior,
(25) $p(\bm{z})=\mathcal{N}(\bm{z};\bm{\mu}_{z},\Sigma_{z})~{}.$
Furthermore, take $\bm{\mu}_{x}=f(W_{x}\bm{z})$ and $\bm{\mu}_{z}=f(W_{z}\bm{z}_{p})$ with $f,W_{x},W_{z}$ defined as before, and $\bm{z}_{p}$ a parameter for the prior (either fixed or learnable). $\Sigma_{x}\in\mathbb{R}^{n_{x}^{2}},\Sigma_{z}\in\mathbb{R}^{n_{z}^{2}}$ are covariance matrices. Thus, $\theta=\{W_{x}\,,W_{z},\,\Sigma_{x},\Sigma_{z}\}$ are the parameters that define the generative model. For convenience, we also define the errors $\bm{\epsilon}_{x}=\bm{x}-\bm{\mu}_{x}$ and $\bm{\epsilon}_{z}=\bm{z}-\bm{\mu}_{z}$ as in section 2. The model is illustrated schematically in fig. 11.
Figure 11. Directed graphical model for PCNs of Rao & Ballard (rao_predictive_1999, ), with variances set to unity. Probabilistic graphical models represent how a joint probability distribution of random variables (nodes) can be factorized, here showing (23) with $\theta={W^{z},W^{z}}$ representing deterministic parameters. The graphical model is surrounded by a plate labelled with $N$ indicating that there are $N$ nodes of this kind, one for each datapoint (see (bishop_pattern_2006, ) for more on graphical models).
@@
$=\frac{1}{2}\bigg{(}(\bm{\epsilon}_{x})^{T}(\Sigma^{x})^{-1}\bm{\epsilon}_{x}+(\bm{\epsilon}_{z})^{T}(\Sigma^{z})^{-1}\bm{\epsilon}_{z}+\ln\left[\det\Sigma_{x}\det\Sigma_{z}\right]\bigg{)}+\text{const.}~{},$. ,
where on the second line we have inserted the Gaussian ansätze (24), (25), and the constant – which is irrelevant to the optimization problem – is $\frac{1}{2}(n_{x}+n_{z})\ln 2\pi$. Choosing diagonal bases so that $\Sigma=\sigma^{2}I$, and disregarding the constant and the factor of 1/2, we have
@@
which is the energy function used by Rao & Ballard (rao_predictive_1999, ) (with variances taken as constants). With a model for $E$ in hand, we can now apply EM, cf. (A.9) and (A.10).

### v2

@@ 2.5.2. Unsupervised Learning
For traditional ANNs, an FNN may be adapted for unsupervised learning by using a special autoencoder architecture, encoding input data into a lower-dimensional representation followed by a decoder which reconstructs the original input. PCNs do not require a special architecture of this sort, but only change the direction of prediction to $\bm{\mu}^{\ell}=f(\bm{w}^{\ell+1}\bm{a}^{\ell+1})$ i.e. the use of a generative PCN. Indeed, PCNs were originally conceived in this way (Rao and Ballard, 1999) (also see section 3). The training procedure for unsupervised learning requires only a minor change to the approach described above. Simply keep the final layer $\bm{a}^{L}$ unclamped during training, keeping data $\bm{x}^{(n)}$ clamped to $\bm{a}^{0}$, and run IL. The model functions like an autoencoder: higher layers take the role of the latent space, which during training encodes increasingly abstract features of the data clamped to the lowest layer. As such, the PCN takes the role of both encoder and decoder.
As for the ‘testing’ procedure, one can sample the latent space as a decoder using a noise vector and ancestral sampling. Seeing the PCNs as a hierarchical probabilistic model (cf. section 3), each layer has a (Gaussian) conditional probability distribution. The idea of ancestral sampling is to generate a sample from the root variable(s) (the top layer) and then sample from the subsequent conditional distributions based on this (Prince, 2023). Finally, a synthetic datapoint at the bottom is obtained (Ororbia and Kifer, 2022).
@@ 2.5.3. Hierarchical PCNs
Discriminative and generative PCNs can together be called hierarchical PCNs, defined by the local prediction $\bm{\mu}^{\ell}=f(\bm{w}^{\ell\pm 1}\bm{a}^{\ell\pm 1})$. The supervised and unsupervised learning that can be done with generative PCNs can then be called the training modes of PCNs. These are illustrated in fig. 7. Intuitively, we can see the difference between discriminative and generative PCNs in the direction of predictions: in the discriminative model, predictions flow from data to labels, while errors flow from labels to data. In generative models, this is reversed. [^10] It should be noted that testing can also be ‘reversed’ in hierarchical PCNs, which we call backwards testing. If one has trained the discriminative PCN, one can clamp labels and find the minimum energy with iterative inference, producing a synthetic image at the bottom. Conversely, a trained generative PCN can classify images by iterative inference, using a clamped image instead of a clamped label. Thus, confusingly, discriminative PCNs can be used for generative tasks, and generative PCNs can be used for discriminative tasks. However, in practice, neither works better than their forward-tested counterpart (Kinghorn et al., 2023; Millidge et al., 2022a), meaning generative PCN are the natural choice for generation, as discriminative PCNs are for classification. This justifies the naming convention used here and elsewhere in the literature.
Figure 7. Overview of hierarchical PCNs. Model type (discriminative vs. generative), learning mode (supervised vs. unsupervised), and direction of predictions and errors are shown. Cf. section 2.2 for an explanation of our convention. For clarity, one neuron per layer is shown, and bias has been left out.
@@ 2.5.4. Empirical Results
Compared to discriminative PCNs, generative PCNs remain underexplored the literature. This is perhaps surprising considering the original conception of PC as an unsupervised generative model. At the same time, as was mentioned, metrics for what counts as a ‘good’ generative model are wide and varied (Prince, 2023), making it a more involved field of study in general. Indeed, the works in the literature are disparate and focus on different aspects.
Although not used for machine learning tasks, we mention (Rao and Ballard, 1999) here as the first generative PCN trained in unsupervised mode. Then, (Salvatori et al., 2021a) used generative PCN as an associative memory model. It was tested on the task of reconstructing from corrupted and partial data, using up to 500 datapoints. They find good performance, outperformining autoencoders, and both Hopfield networks (Hopfield, 1982) and modern Hopfield networks (Krotov and Hopfield, 2016) in most cases. Another set of unsupervised tasks was considered by (Ororbia and Kifer, 2022), and extended in (Ororbia and Mali, 2023). This work used a set of models dubbed neural generative coding (NGC). This framework is equal in spirit to PCNs as presented in section 2, but differs in notation/terminology as well as some structural aspects. The authors study the network’s performance on reconstruction & sampling ability (as measured by the likelihood) on black and white images. They find that NGC is competitive with VAEs and GANs on these tasks, and works well on downstream classification.
Next, a smaller study (Kinghorn et al., 2023) considers the performance of generative PCNs on standard classification tasks, training the model in supervised mode with backwards testing. Similar to (Alonso et al., 2024), it is observed that the weights of different layers are updated at different rates, which causes model accuracy to worsen after it has peaked. They propose that regularizing weights remedies this pathology.
The recent work of (Zahid et al., 2023b) considers the sampling ability of generative PCNs trained in supervised mode with a modified training objective (cf. section A.2.3). They find that their objective improves both likelihood and sample quality, but suffers from increased computation time with results not matching performance of VAE counterparts. An improvement was presented in (Zahid et al., 2024), where the inference phase is combined with Langevin sampling, giving results that match or exceed VAE performance on metrics such as FID (Fréchet Inception Distance), diversity, and coverage.
@@ 2.6. PC Graphs
The previous section showed how the structure of discriminative PCNs, could be extended by changing the definition of local prediction from $\bm{\mu}^{\ell}=f(\bm{w}^{\ell-1}\bm{a}^{\ell-1})$ to $\bm{\mu}^{\ell}=f(\bm{w}^{\ell+1}\bm{a}^{\ell+1})$. This can be taken one step further: PCNs can be naturally generalized to arbitrary graphs, called PC graphs by (Salvatori et al., 2022a). [^11] Note that PC graphs are distinct from graph neural networks (GNNs), which are used for graph-structured data (Ward et al., 2022). These are trained using IL, but dispense with the hierarchical structure of layers. These networks can be understood as a superset of both discriminative and generative PCNs, and can flexibly be used for a variety of tasks.
@@ 2.6.1. Structure
where $w_{ij}\in\mathbb{R}$, and the sum is over all the other nodes, meaning self-connections are left out ($w_{ii}=0$), as in (Salvatori et al., 2022a). This defines a fully-connected PC graph. If we choose to include self-connections one simply takes the sum over all values of $i$. In vector notation, one can write $\bm{\mu}=f(\bm{wa})$.
A small PC graph is illustrated in fig. 8, where each activation node $a_{i}$ and error node $\epsilon_{i}$ has been grouped in a vertex $v_{i}$. Importantly, graphs with different connectivity/topology can be obtained by multiplying the weight matrix with a mask or adjacency matrix. In this way, one obtains the architectures discussed in earlier sections: the discriminative PCN, as well as the generative PCN, depending on which weights are masked, cf. fig. 9.
@@
Figure 9. Mask/adjacency matrices for a 4-node PC graph. Different matrices lead to different architectures, like the discriminative PCN, generative PCN, or an arbitrary topology. Larger networks are obtained in the same way. See also fig. 4 in (Salvatori et al., 2022a).
@@ 2.6.2. Training Procedure
Inference $\Delta\bm{a}^{\ell}=-\gamma\left(\bm{\epsilon}^{\ell}-\left(\bm{w}^{\ell}\right)^{T}\bm{\epsilon}^{\ell\mp 1}\odot f^{\prime}\left(\bm{w}^{\ell}\bm{a}^{\ell}\right)\right)$ $\Delta\bm{a}=-\gamma\left(\bm{\epsilon}-\bm{w}^{T}\bm{\epsilon}\odot f^{\prime}\left(\bm{w}\bm{a}\right)\right)$
@@ 2.6.4. Empirical Results
PC graphs have only been used in (Salvatori et al., 2022a), with models trained using supervised mode only. They train models for classification, generation, reconstruction, denoising, and associative memory tasks (using different adjacency matrices). An interesting result is that for classification using a fully connected model, PC graphs perform much better than other fully connected architectures such as Boltzmann machines and Hopfield networks, up to 30% better on MNIST. At the same time, the fully connected model does not perform comparably to hierarchical networks trained with either BP or IL. This might be expected, considering that depth is suggested to be critical for performance in neural network learning (Prince, 2023). At the same time, such statements are based mostly on empirical evidence, and lack a definitive theoretical basis. For the other tasks considered, only proofs of concepts are provided to demonstrate their flexibility. Unsupervised learning is also not considered.
@@ 2.7. Generalized ANNs
Thus, it becomes clear that formally, PCNs and PCGs can be seen as types of ‘generalized ANNs’ that go beyond hierarchical structures, by virtue of the use of IL as opposed to BP. This is very interesting, for at least two reasons. First, as observed by (Salvatori et al., 2022a), PC graphs allow one to train non-hierarchical structures with a brain-like topology. Speculatively, such networks could, if better understood, share some of the advantages that biological brains have over ANNs, such as vastly superior energy efficiency and parallelization. Second, from a very general perspective, topological considerations have strongly contributed to several advances in machine learning in the past. A prominent example is that of residual networks and skip connections (Zagoruyko and Komodakis, 2016), which have allowed training of much deeper networks, and can improve performance on a variety of tasks (Prince, 2023). In a sense, the very notion of depth in deep learning is a topological feature. As such, we consider this an important avenue for further work.
@@ 3. PCNs as Probabilistic Latent Variable Models
Section 2 introduced the PCN as a type of generalized ANN. In this section we discuss a second, complementary, perspective: PCN as a probabilistic latent variable model. Conceptualizing PCNs in this way allows a principled derivation of equations in the previous section, and brings to light connections to other methods well-known in machine learning (e.g., variational autoencoders and linear factor models). It also reveals a number of assumptions and modeling choices, providing a basis for improvements and possible future developments.
We start our step-by-step derivation from maximum likelihood estimation in probabilistic (Bayesian) models, discussing how expectation maximization (EM) implements maximum likelihood for models with latent variables, and how different generative models lead to PC as formulated by (Rao and Ballard, 1999) or to multi-layer PC as described in the previous section.
{mdframed}
[backgroundcolor=lightergray]
@@ A Brief History of PCNs
The original model of Rao & Ballard (Rao and Ballard, 1999) was conceived as a probabilistic hierarchical model with Gaussian relations between layers. This formulation contained the ingredients typically associated with PC, chief among them the minimization of prediction errors in a hierarchical fashion. Work by Friston (Friston, 2003; Friston et al., 2006; Friston, 2008) showed how the computational steps in PC (minimization of the energy with respect to activations and weights) are an instance of EM, also understood as a variational inference procedure using a variational free energy (i.e., the evidence lower bound) as the objective function. This perspective explains why EM maximizes the likelihood of the data, and can be used to derive generalized objective functions (cf. appendix A).
@@ 3.1. Expectation Maximization
(17) $\operatorname{NLL}(\theta)=-\ln p_{\theta}(\bm{x}),$
where $\theta$ are the parameters that define the model. For some joint distribution $p_{\theta}(\bm{x},\bm{z})$, also called the generative model, it can be shown that the following two steps will minimize the NLL (see appendix A):
@@
[…]
Here, $\ln p_{\theta}(\bm{x},\bm{z})$ is also called complete data log-likelihood, since it represents the likelihood function calculated using the full set of observed and latent variables – to distinguish it from the likelihood calculated only with observed data (Bishop, 2006). We will label the negative complete data log-likelihood as $E$:
@@
which refers to energy, as in section 2 (cf. section 3.4.3 for discussion on energy-based models). We will choose to minimize $E$ as opposed to maximizing $\ln p_{\theta}(\bm{x},\bm{z})$. It has furthermore been shown (cf. appendix A, ref. (Neal and Hinton, 1998)) that partial E-steps and M-steps will also minimize the NLL:
[…]
This, with a change of notation, is incremental IL as presented in section 2. To obtain standard IL, one employs a full E-step, followed by a partial M-step.
@@ 3.2.1. PCNs of Rao & Ballard
and likewise for the prior, $p(\bm{z})=\mathcal{N}(\bm{z};\bm{\mu}_{z},\Sigma_{z})~{}$. Furthermore, take $\bm{\mu}_{x}=f(W_{x}\bm{z})$ and $\bm{\mu}_{z}=f(W_{z}\bm{z}_{p})$ with $f,W_{x},W_{z}$ defined as in section 2, and $\bm{z}_{p}$ a parameter for the prior (either fixed or learnable). $\Sigma_{x}\in\mathbb{R}^{n_{x}^{2}},\Sigma_{z}\in\mathbb{R}^{n_{z}^{2}}$ are covariance matrices. Thus, $\theta=\{W_{x}\,,W_{z},\,\Sigma_{x},\Sigma_{z}\}$ are the parameters that define the generative model. For convenience, we also define the errors $\bm{\epsilon}_{x}=\bm{x}-\bm{\mu}_{x}$ and $\bm{\epsilon}_{z}=\bm{z}-\bm{\mu}_{z}$ as in section 2. The model is illustrated schematically in fig. 11.
Figure 11. Directed graphical model for PCNs of Rao & Ballard (Rao and Ballard, 1999), with variances set to unity. Probabilistic graphical models represent how a joint probability distribution of random variables (nodes) can be factorized, here showing (23) with $\theta=\{W_{z},W_{z}\}$ representing deterministic parameters. The graphical model is surrounded by a plate labelled with $N$ indicating that there are $N$ nodes of this kind, one for each datapoint (cf. e.g.(Bishop, 2006) ).
@@
$=\frac{1}{2}\bigg{(}(\bm{\epsilon}_{x})^{T}(\Sigma_{x})^{-1}\bm{\epsilon}_{x}+(\bm{\epsilon}_{z})^{T}(\Sigma_{z})^{-1}\bm{\epsilon}_{z}+\ln\left[\det\Sigma_{x}\det\Sigma_{z}\right]\bigg{)}+\text{const.}~{},$. ,
where on the second line we have inserted the Gaussian ansätze (24), the prior, and the constant – which is irrelevant to the optimization problem – is $\frac{1}{2}(n_{x}+n_{z})\ln 2\pi$. Choosing diagonal bases so that $\Sigma=\sigma^{2}I$, and disregarding the constant and the factor of 1/2, we have
@@
which is the energy function used by Rao & Ballard (Rao and Ballard, 1999) (with variances taken as constants). With a model for $E$ in hand, we can now apply EM, cf. (18) and (19).